# Cosmological Inflation and Dark Sector from 11D Supergravity

### v2

@@ 5 Conclusions and Discussion
In this paper, we have discussed multiple potential methods for compactifying 11D supergravity, which can be represented as $\mathcal{M}^{3,1}\times T_{q_{1}}^{2}\times T_{q_{2}}^{2}\times T_{q_{3}}^{2}\times S^{1}$. As a low-energy effective theory derived from the ultraviolet-complete M-theory, the inherent completeness of quantum gravity in the ultraviolet regime does not forbid the existence of dark matter and dark energy in a four-dimensional FRW universe. The dynamics of the inflaton, Standard Model of Particle Physics, dark matter, and dark energy can be all included in the effective action in four-dimensional spacetime. Meanwhile, the inflaton can be interpreted as the conformal vibrations of extra dimensions.
@@
In our examples of driven potential (71), the dynamics of sub-Planck scale extra dimensions govern the cosmological inflation that occurs in four-dimensional spacetime. When the scale function has minimum value $\Phi_{min}=-\frac{1}{n}$ and $s_{1},s_{2},s_{3}$ take values $s_{1}=\frac{1}{2},s_{2}=0,s_{3}\geq 1$, the theoretical predictions align with the observational range derived from the 2021 BICEP/Keck results, see Figure 2. The special case addressed in Subsection 2.1 satisfies all the observational constraints from both 2021 BICEP/Keck and 2018 Planck results. In this scenario, the geometric configuration of the extra dimensions is $S^{2}\times T_{2}^{2}\times T^{2}_{2}\times S^{1}$. As inflation progresses, the three tori undergo a contraction in size. Once the inflaton $\phi$ initiates its decay into gravitino $\psi_{A}$ and 3-form $A_{MNP}$, inflation exits. During this decay process, particles encompassed by the Standard Model of Particle Physics, as well as dark matter and dark energy, are generated. Then the extra dimensions become inactive and stabilize into a fixed shape, where the standard sizes of the extra dimensions are around $10^{5}l_{p}$.
@@
As mentioned in Section 4, the quantum effects from extra dimensions are detectable in principle through high-energy experiments and astronomy observations. However, detecting the inflaton in accelerators is extremely challenging in practice. Nevertheless, there are other possible ways to indirectly infer the existence of extra dimensions, such as observations of the power spectrum of Cosmic Microwave Background (CMB) and the rotation curve of galaxies. On the one hand, in our compactification methods, the supersymmetry of gravity in four-dimensional spacetime is spontaneous breaking. This theoretical prediction aligns with experimental observations, as there is currently no conclusive evidence for superpartners of the graviton and the particles in the Standard Model of Particle Physics. On the other hand, the geometric configuration of the extra dimensions is reflected by the inflaton field, which can be interpreted as a component of dark matter or dark energy in the late universe. The modified Newton’s gravity law, taking into account a spherically symmetric distribution of the inflaton, is presented in (113)(118)(132). The equation (118), which depicts a point particle surrounded by dark matter, can elucidate the Tully-Fisher relation given in (124). The range of the parameter $m_{2}$ in (118) can be constrained through observations for galaxies. It is worth mentioning that our method to derive the properties of the dark sector is fixed on the Minkowski background and the backreaction sourced by matter to spacetime is neglected since the gravitational field is very weak. To study the case in the strong gravitational field, more precise calculations regarding the backreaction should be performed, which is left for future work.
APPENDIX
@@ ACKNOWLEDGEMENTS
We would like to thank Chi-Ming Chang, Zhengyuan Du, Wen-Xin Lai, Wei Song, Yi-xiao Tao, Dongjian Wu, and Tianqing Zhu for useful discussions. The work is partially supported by the national key research and development program of China NO. 2020YFA0713000.

### v3

@@ V Conclusions and Discussion
In this paper, we have discussed multiple potential methods for compactifying 11D supergravity, which can be represented as $\mathcal{M}^{3,1}\times T_{q_{1}}^{2}\times T_{q_{2}}^{2}\times T_{q_{3}}^{2}\times S^{1}$. As a low-energy effective theory derived from the ultraviolet-complete M-theory, the inherent completeness of quantum gravity in the ultraviolet regime does not forbid the existence of dark matter and dark energy in an FRW universe with four extended spacetime dimensions and seven extremely small compactified spatial dimensions. The dynamics of the inflaton, Standard Model of Particle Physics, dark matter, and dark energy can be all included in the effective action for the four-dimensional extended spacetime. Meanwhile, the inflaton can be interpreted as the conformal vibrations of extra dimensions.
@@
In our examples of driven potential (45), the dynamics of sub-Planck scale extra dimensions govern the cosmological inflation that occurs in four-dimensional extended spacetime. When the scale function has minimum value $\Phi_{min}=-\frac{1}{n}$ and $s_{1},s_{2},s_{3}$ take values $s_{1}=\frac{1}{2},s_{2}=0,s_{3}\geq 1$, the theoretical predictions align with the observational range derived from the 2021 BICEP/Keck results, see Figure 2. The special case addressed in Subsection II.1 satisfies all the observational constraints from both 2021 BICEP/Keck and 2018 Planck results. In this scenario, the geometric configuration of the extra dimensions is $S^{2}\times T_{2}^{2}\times T^{2}_{2}\times S^{1}$. As inflation progresses, the three tori undergo a contraction in size. Once the inflaton $\phi$ initiates its decay into gravitino $\psi_{A}$ and 3-form $A_{MNP}$, inflation exits. During this decay process, particles encompassed by the Standard Model of Particle Physics, as well as dark matter and dark energy, are generated. Then the extra dimensions become inactive and stabilize into a fixed size, where the standard sizes of the extra dimensions are around $10^{5}l_{p}$.
@@
As mentioned in Section IV, the quantum effects from extra dimensions are detectable in principle through high-energy experiments and astronomy observations. However, detecting the inflaton in accelerators is extremely challenging in practice. Nevertheless, there are other possible ways to indirectly infer the existence of extra dimensions, such as observations of the power spectrum of Cosmic Microwave Background (CMB) and the rotation curve of galaxies. On the one hand, in our compactification methods, the supersymmetry of gravity in four-dimensional extended spacetime is spontaneous breaking. This theoretical prediction aligns with experimental observations, as there is currently no conclusive evidence for superpartners of the graviton and the particles in the Standard Model of Particle Physics. On the other hand, the geometric configuration of the extra dimensions is reflected by the inflaton field, which can be interpreted as a component of dark matter or dark energy in the late universe. The modified Newton’s gravity law, taking into account a spherically symmetric distribution of the inflaton, is presented in (86)(91)(105). The equation (91), which depicts a point particle surrounded by dark matter, can elucidate the Tully-Fisher relation given in (97). The range of the parameter $m_{2}$ in (91) can be constrained through observations for galaxies. It is worth mentioning that our method to derive the properties of the dark sector is fixed on the Minkowski background and the backreaction sourced by matter to spacetime is neglected since the gravitational field is very weak. To study the case in the strong gravitational field, more precise calculations regarding the backreaction should be performed, which is left for future work.
Appendix A Gauge Conditions
Because the gauge parameter $\theta_{MN}$ has 55 degrees of freedom, we can impose gauge conditions as follows:
1. For a given function symmetric tensor $\mathcal{T}_{\mu\nu}^{\text{(matter)}}[x^{\mu}]$, the equations
$T_{\mu\nu}^{(A,\psi)}[x^{A}]=\frac{m_{p}^{2}}{m_{p(11)}^{9}}\mathcal{T_{\mu\nu}^{\text{(matter)}}}[x^{\mu}],$ (106)
expect for $\mu=\nu=0$ case, restrict 9 degrees of freedom.
$G_{a_{i}A}$ $=0,(A\neq a_{i}\text{ and }i=1,2,3)$ (107)
$G_{\mu A}$ $=0,(A\neq\mu)$ (108)
$G_{10,A}$ $=0,(A\neq 10)$ (109)
where the operator $\delta(\cdots)$ is the variation with respect to supersymmetry. The supersymmetric transformation law for the metric $G_{AB}$ can be expressed as
$\delta G_{a_{1}b_{1}}$ $=g_{a_{1}b_{1}}\delta\Phi,$ (118)
$\delta G_{a_{2}b_{2}}$ $=0,$ (119)
$\delta G_{a_{3}b_{3}}$ $=2g_{a_{3}b_{3}}\Phi\delta\Phi,$ (120)
$\delta G_{10,10}$ $=-6\Phi^{-7}\delta\Phi.$ (121)
The gauge condition (117) indicates the absence of gravitino, $\psi_{\mu}=0$, in the four-dimensional extended spacetime. Consequently, supersymmetry of gravity in the four-dimensional extended spacetime is spontaneous breaking. As there is no strong observable signal from supersymmetric field theory in 4D, we postulate that $\Phi$ has no superpartner, thus assuming $\delta\Phi=0$. Hence, the compactification methods spontaneously break all the $\mathcal{N}=32$ supersymmetries present in the eleven-dimensional bulk, where the classical solutions of (117)(118)(119)(120)(121) are $\psi_{M}=0$. The non-trivial solution to the gauge parameter $\epsilon$ can be derived from the equation
Note that there are trivial solutions $\epsilon=0$, when the 3-form takes trivial solution $A_{MNP}=0$. After imposing the gauge conditions (106)(107)(108) and (109)(115)(122), the only two remaining undetermined functions are the scale factor $a(t)$ and the scale function $\Phi(x^{\mu})$ in (15).
In the four-dimensional extended spacetime, we anticipate that Einstein’s gravity predominantly governs the laws of classical gravity as seen in (16). Moreover, gauge conditions (106) imply that it is possible to obtain a four-dimensional effective action, in which the solutions to the equations of motion are represented by $\Phi(x^{\mu})$ and $a(t)$. While the equations of (106), expect for $\mu=\nu=0$, have already been utilized as gauge conditions to determine the solution of 3-form, these equations do not establish the relation between $\Phi(x^{\mu})$ and $a(t)$. To clarify the relationship between $\Phi(x^{\mu})$ and $a(t)$, we introduce an additional equation, which is the same as (106) but only for the case $\mu=\nu=0$. Subsequently, the solution for $\mathcal{T}^{\text{(matter)}}_{00}[\psi_{A},A_{MNP},a(t),\Phi(x^{\mu}),x^{\mu}]$ obtained from this equation uniquely defines the Lorentz-invariant symmetric tensor $\mathcal{T}^{\text{(matter)}}_{\mu\nu}[\psi_{A},A_{MNP},a(t),\Phi(x^{\mu}),x^{\mu}]$, thereby imposing 9 gauge conditions specified in (106). These 9 gauge conditions, along with (107)(108)(109)(115)(122), constrain the solutions $\psi_{A}(a(t),\Phi(x^{\mu}),x^{A})$ and $A_{MNP}(a(t),\Phi(x^{\mu}),x^{A})$ in such a way that they determine $\mathcal{T}^{\text{(matter)}}_{\mu\nu}[a(t),\Phi(x^{\mu}),x^{\mu}]$ uniquely. This tensor can be derived by the equations of motion for $\Phi(x^{\mu})$ and $a(t)$ in four-dimensional extended spacetime. The equations (106) for $\mu,\nu=0,1,2,3$ serves to define an effective energy-stress tensor $\mathcal{T}_{\mu\nu}$ as
where $\mathcal{T}_{\mu\nu}^{\text{(matter)}}$ is the stress tensor of some four-dimensional matter. Hence, in the four-dimensional extended spacetime, the equation of motion (17) implies 4D effective Einstein’s equation
$\mathcal{G}_{\mu\nu}=\frac{1}{m_{p}^{2}}\mathcal{T}_{\mu\nu},$ (124)
and the associated effective action is given by
where
The gravitino field $\psi_{A\dot{\alpha}\dot{\alpha_{1}}\dot{\alpha_{2}}\dot{\alpha_{3}}}$ gives rise to 56 fermion fields $\psi_{i\dot{\alpha}\dot{\alpha}_{1}\dot{\alpha_{2}}\dot{\alpha_{3}}}$ for $i=4,5,\cdots,10$ in the 4D extended spacetime, which contains the fermion fields in the Standard Model of Particle Physics. Additionally, there are 21 spin-1 boson fields $A_{\mu ij}$, 35 spin-0 scalar fields $A_{ijk}$, and 7 spin-0 tensor fields $A_{\mu\nu k}$ for $i,j,k=4,5,\cdots,10$. These fields contain ordinary matter, dark matter, and dark energy. Then, the lagrangian for matter in the 4D extended spacetime can encompass ordinary matter from the Standard Model of Particle Physics $\mathcal{L}_{\text{SM}}$, as well as dark matter $\mathcal{L}_{\text{DM}}$ and dark energy $\mathcal{L}_{\text{DE}}$,
$\mathcal{L}_{\text{matter}}=\mathcal{L}_{\text{SM}}+\mathcal{L}_{\text{DM}}+\mathcal{L}_{\text{DE}},$ (127)
and the associated effective action for the 4D extended spacetime is (125). This allows us to derive a cosmological inflation model based on a specified driven potential $V(\Phi)$, where the potential can be derived by integrating out the coordinates dependence of extra dimensions from the 11D supergravity action (2).
Appendix B Properties of Potential
@@
Appendix C Exit Inflation
There exist additional matter terms within the four-dimensional effective action of 11D Supergravity (125). Because the 11D metric $G_{AB}$ couples to gravitino $\psi_{A}$ and 3-form $A_{MNP}$, the inflaton $\phi$, associated to the size function $\Phi$ of extra dimensions, will couple to some components of $\psi_{A}$ and $A_{MNP}$ after compactification. When the inflaton $\phi$ begins to decay into the gravitino and the 3-form, the inflation exits. In this period, we keep the supersymmetry gauge condition $\delta g_{\mu\nu}=0$ and release $\delta g_{a_{1}b_{1}},\delta h_{a_{2}b_{2}},\delta\gamma_{a_{3}b_{3}}\neq 0$, see A for more details about the supersymmetry gauge condition during inflation period. This decay process gives rise to a non-zero $\mathcal{L}_{\text{matter}}$ term in the four-dimensional effective action. For example, matter in the universe can be generated by introducing the interactions between the inflaton and the Higgs field $H$, expressed as $\xi\phi^{2}H^{2}$, into the Lagrangian $\mathcal{L}_{\text{matter}}$.
@@ ACKNOWLEDGEMENTS
We would like to thank Chi-Ming Chang, Zhengyuan Du, Wen-Xin Lai, Jian Xin Lu, Wei Song, Yi-xiao Tao, Dongjian Wu, and Tianqing Zhu for useful discussions. The work is partially supported by the national key research and development program of China NO. 2020YFA0713000.